OpenStudy (anonymous):

if 2 prisms have a side ratio of 3:4, and the smaller one has a surface area of 300sq in what is the SA of the larger?

hero (hero):

In general, two similar solids have the following properties: Length ratio - a:b Area ratio - a^2:b^2 Volume ratio - a^3:b^3 In this case, we have two similar prisms with: Length ratio - 3:4 , which means Area ratio - 3^2: 4^2 If the surface area of the smaller prism is 300 sq ft, then this means we setup the following ratio: 300/9 = x/16 In other words, 9x = 300*16 9x = 4800 x = 4800/9 x = 533.333 Therefore, the larger prism has a surface area of 533 and 1/3 sq. ft

hero (hero):

btw, ikuudkme

OpenStudy (anonymous):

haha cool thanks